# A Practical Examination of AI-Generated Text Detectors for Large Language Models

### v3

@@ 4 Experiment
Table 3: Performance of different detectors across the entire dataset
Text Generation and Detection Process. Once the prompt samples were selected, we needed to generate positive examples. The process for this can be seen in Figure 1. We employ three different strategies for prompting the models. The first strategy involves using a basic prompt for each domain that explains the goal of the model and the desired output format. The second strategy consists of requesting that the model be as human as possible. The third strategy requests that the model rewrite and improve upon the human written response  [^6] Prompts and templates in appendix. The first strategy aims to simulate a basic system prompt that would generally be in place on a model someone is using to generate content. The second strategy simulates the case where a user might try to get the model to generate content that closely resembles human-generated content. The third strategy simulates a scenario where the user writes their own response and simply wants the model to clean it up or make it easier to understand. The outputs of the models were taken as is with no editing. After generating the positive examples, we passed all of the machine-generated and human-generated examples through the detectors. RADAR, Fastdetectgpt, Wild, and T5Sentinel all return a percentage probability for each class, and GPTID and LogRank return a value representing their score. We do not use any thresholds and take the scores as is for AUROC and TPR@FPR metrics.
@@ 5 Results and Analysis
Figure 4: Correlations between various FPR rates and the overall AUROC score. AUROC score is much more representative of the middle FPR rates, while this detection task is much more concerned with the lower end of FPR.
@@ 5.2 Template Prompting
Code Reviews Abstract QA Summ Dialogue Trans. Arena Score
Model TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC
@@ 5.4 TPR@FPR vs AUROC
In this paper, we utilize both the AUROC and TPR@FPR metrics. However, we also argue that TPR at a low FPR is a much more important metric for this detection task. Figure 4 shows the correlation between TRP scores at various FPR rates and the AUROC score for all tasks, detectors, and models used in this research. The AUROC correlates much higher with FPR rates in the 0.4 to 0.6 range and much lower with FPR rates at the edges, less than 0.2 and greater than 0.8. While the 0.75 is still a reasonable correlation value, the AUROC is still much more representative of the middle FPR’s while we are really concerned with the lower FPR’s for this task. This is why we report the TPR@0.01, which is much more representative of the applicability of a detector than the AUROC.
@@ Appendix A More Results
This section contains results for detections by models and tasks. It also includes the prompts used in plain prompting.
@@ A.2 Plain Prompts
Table 12 shows the prompts used for each task in the plain prompting.
@@
Question Answering You are a helfpul question answering assistant that will answer a single quesetion as completely as possible given the information in the question. Do NOT using any markdown, bullet, or numbered list formatting. The assistant will use ONLY paragraph formatting. **Respond only in {language}**
Summarization You are a helfpul summarization assistant that will summarize a given article. Provide only the summarization in paragraph formatting. Do not introduce the summary. **Respond in {language}**
@@
Rewriting You are a helpful writting assistant. Rewrite the following text to improve clarity and professionalism. Do not provide any other text. Only provide the rewritten text.

### v4

@@ 4.2 Text Generation and Detection Process
Table 4: Performance of different detectors across the entire dataset.
@@ 5.2 Adversarial Prompting
Model Code Reviews Abstract QA Summ Dialogue Trans. Arena Score
TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC
@@ 5.3 Rewriting
Figure 4: Correlations between the TPR at various FPR rates and the overall AUROC score. AUROC score is more representative of the middle FPR rates, while this detection task is more concerned with the lower end of FPR.
@@ 5.4 TPR@FPR vs AUROC
In this paper, we utilize both the AUROC and TPR@FPR metrics. However, we also argue that TPR at a low FPR is a much more important metric for this detection task. Figure 4 shows the correlation between TPR scores at various FPR rates and the AUROC score for all tasks, detectors, and models used in this research. The AUROC correlates much higher with FPR rates in the 0.4 to 0.6 range and much lower with FPR rates at the edges, less than 0.2 and greater than 0.8. While the 0.75 is still a reasonable correlation value, the AUROC is still much more representative of the middle FPR’s while we are really concerned with the lower FPR’s for this task. This is why we report the TPR@0.01, which is much more representative of the applicability of a detector than the AUROC.
@@ 5.5 Output Quality and Detection
Figure 5: AUROC and TPR@0.01 for each zero-shot method using various underlying models. Only Fast-DetectGPT and Binoculars show a significant change in performance with GPT2-Medium.
5.6 Impact of Model on Zero-shot Methods
Each zero-shot method used in this paper has an underlying model that assists in the detection process. In this paper we consider the model chosen by the respective authors of each detector to be a part of the detector itself. However, we also swapped out each model to directly compare the statistical methods themselves, removing any impact from a specific model on a detector.
Figure 5 shows the results of running each detector across the entire dataset with three different models. There is generally not much of a difference in the ability of a detection method when changing the underlying model. Fast-DetectGPT and Binoculars show a small change in AUROC and a larger change in TPR@0.01 when using the gpt2-medium model (Radford et al., 2019). Gpt2-medium is the oldest model of the three, which likely results in its output logits being different than the generation models more often. This provides some evidence that these zero-shot methods will require updated underlying models to remain successful on more advanced generation models, but more research would need to be conducted.
@@ 6 Conclusion
The detection results for both the Translation task and Rewrite prompt are generally lower than the average detectability for other machine generated text. This encourages a discussion about whether this type of text should be detected as machine generated or not. The text may have been machine generated, but it is heavily influenced by human generated text. Specifically in the translation case, the text should match the source text in another language. In the Rewrite case, the model is not generating any new ideas, just improving the readability of the text. It is clear from the results that these cases are harder to detect than when a model has a more open-ended generation. It is likely worth differentiating between these two cases which we leave to future work.
@@ 7 Limitations
Acknowledgements
L.L. is partly supported by a gift from Hydrox AI and a CMU CyLab Seed Fund. The views expressed are those of the author and do not reflect the official policy or position of the funding agencies.
@@ Appendix A More Results
This section contains results for detections by models and tasks, the prompts used for plain prompting, and the results by detector.
@@ A.1 Results by Model
The following tables show the results for each detector by generation model and task. As discussed in the paper, there is not a significant difference in detectability of a text by the model that generated that text. More specifically, a higher quality model like GPT-4o is not noticeably harder to detect than a lower quality model like Phi-3. The differences in detectability are more obvious across tasks than generation models.
@@ A.2 Plain Prompts
Table 13 shows the prompts used for each task in the plain prompting. The prompts we used were intentionally very simple and not overly instructive. This is because we wanted to replicate a realistic scenario of an average person prompting a language model. We performed small ablations on these prompts and found no difference in detectability.
@@
Question Answering You are a helpful question answering assistant who will answer a single question as completely as possible given the information in the question. Do NOT use any markdown, bullet, or numbered list formatting. The assistant will use ONLY paragraph formatting. **Respond only in {language}**
Summarization You are a helpful summarization assistant that will summarize a given article. Provide only the summarization in paragraph formatting. Do not introduce the summary. **Respond in {language}**
@@
Rewriting You are a helpful writing assistant. Rewrite the following text to improve clarity and professionalism. Do not provide any other text. Only provide the rewritten text.
@@ A.3 Results by Detector
This section shows the numerical value of each detector on each task. In the paper we display graphs representing most of these values but show all of the numbers here for reference. The TPR@.01 and AUROC change significantly across tasks for every detector signifying that these detectors are not equally capable of detecting all types of machine generated text.